Two arigles are supplementary. One angle measures 26 angles. Show all your work!! 4 points) s. than tes are supplementary. One angle measures 26e more than the other. Pind the measure of the two
Answer #1

Supplementary angles means that the sum of all the angles given = 180o.

Let one angle A = x, then the other angle B = x + 26.

Since these are supplementary, A + B = x + x + 26 = 180

Therefore 2x = 180 - 26 = 157

x = 154/2 = 77

Angle A = x = 77o.

Angle B = x + 26 = 77 + 26 = 103o.

(we see that 77 + 103 = 180)